# From Lipschitz Embedding to Lipschitz Equivalence between Dust-like Self-similar Sets

### v1

@@ 4. Lipschitz surjections with measure-preserving properties
$=\lambda_{\bm{j}}^{-1}a_{t}\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})\geq\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}^{-}}\Delta_{K}\geq\frac{\Delta_{F}}{L\delta^{n}}\cdot\delta^{n}\cdot\Delta_{K}=L^{-1}\Delta_{F}\Delta_{K}$
@@
Assume that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$. Then
$\Big{(}\sum_{t=1}^{p}\#\mathcal{I}_{t}\Big{)}\cdot(\underline{r}\delta^{n})^{s}\mathcal{H}^{s}(K)$ $\leq\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))$ (recall (4.4))
$\leq\widetilde{c}\mathcal{H}^{s}(F_{\bm{j}})$
$\leq\frac{\mathcal{H}^{s}(K)}{\mathcal{H}^{s}(F)}\cdot\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(F)$ (by Corollary 4.2)
$=\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(K)$
$\leq\frac{(L\delta^{n})^{s}}{\Delta_{F}^{s}}\mathcal{H}^{s}(K),$ (since $\bm{j}\in J_{n}$)
which implies that $\sum_{t=1}^{p}\#\mathcal{I}_{t}\leq\lfloor\frac{L^{s}}{\underline{r}^{s}\Delta_{F}^{s}}\rfloor=P$. So $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$.
Note that the inclusion $\widetilde{f}:\widetilde{f}^{-1}(F_{\bm{j}})\to F_{\bm{j}}$ naturally generates a Lipschitz surjection
@@
which is a convex combination. This identity should be $\widetilde{c}=\sum_{\bm{j}\in J_{n}}\lambda_{\bm{j}}^{s}\widetilde{c}$. Otherwise, since every term in that combination is positive, there exists $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}<\widetilde{c}$, which contradicts Lemma 4.3. In particular, $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$ for every $\bm{j}\in J_{n}$. This in turn implies by Lemma 4.3 that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$. ∎
@@
Write $r$ to be the common ratio of maps in $\Phi$. By Lemma 2.4, it suffices to show that $\frac{\log\lambda_{j}}{\log r}\in\mathbb{Q}$ for $j\in J$. Let $j\in J$ and pick for $n\geq 1$ a positive integer $k_{n}$ so that
@@
Since $r_{i}\equiv r$, there exists $q_{t,n}\in\mathbb{N}$ such that $I^{\prime}_{t,n}\subset I^{q_{t,n}}$, where $I^{\prime}_{t,n}$ was defined in (4.3). Moreover, for $t\neq t^{\prime}$, we always have $a_{t}r^{q_{t,n}}\leq\delta^{n}<a_{t^{\prime}}r^{q_{t^{\prime},n}-1}$ and hence $r^{q_{t^{\prime},n}-q_{t,n}}>a_{t}r/a_{t^{\prime}}$. So there is a constant $C\in\mathbb{N}$ such that
(4.5) $|q_{t,n}-q_{t^{\prime},n}|\leq C,\quad\forall 1\leq t,t^{\prime}\leq m,\forall n\in\mathbb{Z}^{+}.$
By (4.4) and Proposition 4.4,
(4.6) $\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\Big{\{}\mathcal{H}^{s}(K)\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{q_{t,n}s}:c_{n,t}\in\mathbb{N}\text{ with }\sum_{t=1}^{p}c_{n,t}\leq P\Big{\}}.$
Note that every inner sum on the right hand side above satisfies
$\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{q_{t,n}s}$ $=r^{q_{1,n}s}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{(q_{t,n}-q_{1,n})s}$
$\in\Big{\{}r^{q_{1,n}s}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{\beta s}:-C\leq\beta\leq C\text{ is an integer}\Big{\}}.$ (by (4.5))
Combining this with (4.6), we see that
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\{\mathcal{H}^{s}(K)r^{q_{1,n}s}b:b\in\mathcal{B}\},\quad\forall n\geq 1,$
where
$\mathcal{B}:=\Big{\{}\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{\beta s}:c_{n,t}\in\mathbb{N},\beta\in\mathbb{Z}\text{ are such that}\sum_{t=1}^{p}c_{n,t}\leq P\text{ and }-C\leq\beta\leq C\Big{\}}$
is a finite collection. Therefore, we can find (by the pigeonhole principle) a pair of distinct integers $n,\tilde{n}$ such that
@@
Unfortunately, the above argument fails without the homogeneity of $\Phi$, and even granting the homogeneity of $\Psi$ seems to offer little help. This reveals a subtle asymmetry: to upgrade a Lipschitz embedding to a Lipschitz isomorphism, assuming the homogeneity on $K$ is markedly easier than assuming that on $F$. It is noteworthy that Proposition 4.4 may relate to a characterization given by Xi [20] via finding appropriate graph-directed subsystems, but we do not know whether the measure-preserving property would directly ensure that.
@@
    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0$ unless $(t,\bm{i})=(t^{\prime},\bm{i}^{\prime})$;
@@
    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))\gtrsim\mathcal{H}^{s}(F_{\bm{j}})$ for all $1\leq t\leq m$ and $\bm{i}\in\mathcal{I}_{t}$.
@@
For (1), write $E:=\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})$ and suppose on the contrary that $\mathcal{H}^{s}(E)>0$. By Lemma 3.4, we can find $\eta_{k}\in J_{k}$ with $F_{\eta_{k}}\subset F_{\bm{j}}$ such that
(4.7) $\lim_{k\to\infty}\frac{\mathcal{H}^{s}(E\cap F_{\eta_{k}})}{\mathcal{H}^{s}(F_{\eta_{k}})}=1.$
For all large $k$, consider $A_{k}:=\widetilde{f}^{-1}(F_{\eta_{k}})$ and let
@@
Clearly, $A_{k,1}$, $A_{k,2}$ are both non-empty (since $\mathcal{H}^{s}(E\cap F_{\eta_{k}})>0$), mutually disjoint (since $(t,\bm{i})\neq(t^{\prime},\bm{i}^{\prime})$) and $A_{k,1}\cup A_{k,2}\subset A_{k}$. Moreover, it follows from (4.7) that
@@
On the other hand, by Lemma 4.3, we can write $\widetilde{f}^{-1}(F_{\eta_{k}})=\bigcup_{\ell=1}^{p}\bigcup_{\omega\in\mathcal{I}_{\ell,k}}(a_{\ell}O_{\ell}K_{\omega}+b_{\ell})$ for some $\mathcal{I}_{\ell,k}\subset I^{\prime}_{\ell,k}$. So
$A_{k,1}=\bigcup_{\omega\in\mathcal{I}_{t,k}:K_{\omega}\subset K_{\bm{i}}}(a_{t}O_{t}K_{\omega}+b_{t})\quad\text{and}\quad A_{k,2}=\bigcup_{\omega\in\mathcal{I}_{t^{\prime},k}:K_{\omega}\subset K_{\bm{i}^{\prime}}}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}}).$
By Proposition 4.4, $\psi_{\eta_{k}}^{-1}\widetilde{f}^{-1}(F_{\eta_{k}})\in\mathcal{K}$. So $\psi_{\eta_{k}}^{-1}(A_{k,1}),\psi_{\eta_{k}}^{-1}(A_{k,2})\in\mathcal{K}$ (because they are sub-unions). Also,
$\mathcal{H}^{s}(A_{k,2})$ $\geq\min_{\omega\in I^{\prime}_{t^{\prime},k}}\mathcal{H}^{s}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}})$
$\geq(\underline{r}\delta^{k})^{s}\mathcal{H}^{s}(K)$ (recall (4.3))
$\geq\frac{\underline{r}^{s}\Delta_{F}^{s}}{L^{s}}\cdot\lambda_{\eta_{k}}^{s}\mathcal{H}^{s}(K)$ (since $\eta_{k}\in J_{k}$)
@@
$\psi_{\eta_{k}}^{-1}(x)$ $\mapsto\psi_{\eta_{k}}^{-1}\widetilde{f}(x).$
Since $\psi_{\eta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$ and $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$, we can find by Lemma 2.2 a Lipschitz map $g:E\to F$ with $E\in\mathcal{K}$, $\operatorname{Lip}(g)\leq L$ such that
$\mathcal{H}^{s}(E)\leq\sup_{k}\mathcal{H}^{s}(\psi_{\eta_{k}}^{-1}(A_{k,1}))=\mathcal{H}^{s}(F)\cdot\sup_{k}\frac{\mathcal{H}^{s}(A_{k,1})}{\mathcal{H}^{s}(F_{\eta_{k}})}<\widetilde{c}\mathcal{H}^{s}(F)$
@@
$\mathcal{H}^{s}(g(E))$ $\geq\limsup_{k\to\infty}\mathcal{H}^{s}(g_{k}(\psi_{\eta_{k}}^{-1}(A_{k,1})))$
@@
The latter implies that $g(E)=F$. However, by our choice of $\widetilde{K}$ (recall Corollary 4.2),
$\widetilde{c}\mathcal{H}^{s}(F)>\mathcal{H}^{s}(E)\geq\mathcal{H}^{s}(\widetilde{K})=\widetilde{c}\mathcal{H}^{s}(F),$
which is a contradiction. This proves (1).
For (2), let $1\leq t\leq p$, $\bm{i}\in\mathcal{I}_{t}\subset I^{\prime}_{t,n}$ and let
$B:=\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\Big{(}\bigcup_{(\bm{i}^{\prime},t^{\prime})\neq(\bm{i},t)}\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})\Big{)},$
which is a closed set. We have seen in (1) that $\mathcal{H}^{s}(B)=0$. By Remark 4.5, $\mathcal{H}^{s}(\widetilde{f}^{-1}(B))=\widetilde{c}\mathcal{H}^{s}(B)=0$ and
$=\widetilde{c}^{-1}\mathcal{H}^{s}\big{(}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cup\widetilde{f}^{-1}(B)\big{)}$
$=\widetilde{c}^{-1}\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
$\gtrsim\mathcal{H}^{s}(F_{\bm{j}}),$
where the last estimate again comes from (4.3) and Lemma 3.3(1). This proves (2). ∎

### v2

@@ 4. Lipschitz surjections with measure-preserving properties
$=\lambda_{\bm{j}}^{-1}a_{t}\operatorname{dist}(K_{\bm{i}},K_{\bm{i}^{\prime}})$
$\geq\lambda_{\bm{j}}^{-1}a_{t}r_{\bm{i}^{-}}\Delta_{K}\geq\frac{\Delta_{F}}{L\delta^{n}}\cdot\delta^{n}\cdot\Delta_{K}=L^{-1}\Delta_{F}\Delta_{K}=\gamma$
@@
Now assume that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}\leq\widetilde{c}$. Then since $\mathcal{H}^{s}(\widetilde{K})\leq\mathcal{H}^{s}(K)$ (see Corollary 4.2) and $\bm{j}\in J_{n}$,
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))\leq\widetilde{c}\mathcal{H}^{s}(F_{\bm{j}})=\frac{\mathcal{H}^{s}(\widetilde{K})}{\mathcal{H}^{s}(F)}\cdot\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(F)\leq\lambda_{\bm{j}}^{s}\mathcal{H}^{s}(K)\leq\frac{(L\delta^{n})^{s}}{\Delta_{F}^{s}}\mathcal{H}^{s}(K).$
On the other hand, the expression (4.5) immediately implies
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))\geq\Big{(}\sum_{t=1}^{p}\#\mathcal{I}_{t}\Big{)}\cdot(\underline{r}\delta^{n})^{s}\mathcal{H}^{s}(K).$
Thus $\sum_{t=1}^{p}\#\mathcal{I}_{t}\leq\lfloor\frac{L^{s}}{\underline{r}^{s}\Delta_{F}^{s}}\rfloor=P$. Altogether, we have $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$.
Furthermore, the inclusion $\widetilde{f}:\widetilde{f}^{-1}(F_{\bm{j}})\to F_{\bm{j}}$ naturally generates a Lipschitz surjection
@@
which is a convex combination. The identity should be $\widetilde{c}=\sum_{\bm{j}\in J_{n}}\lambda_{\bm{j}}^{s}\widetilde{c}$. Otherwise, since every term in that combination is positive, there exists $\bm{j}\in J_{n}$ such that $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}<\widetilde{c}$, contradicting Lemma 4.3. In particular, $\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}}))}{\mathcal{H}^{s}(F_{\bm{j}})}=\widetilde{c}$ for every $\bm{j}\in J_{n}$, which in turn implies by Lemma 4.3 that $\psi_{\bm{j}}^{-1}\widetilde{f}^{-1}(F_{\bm{j}})\in\mathcal{K}$. ∎
@@
Write $r$ to be the common ratio of similitudes in $\Phi$. By Lemma 2.4, it suffices to show that $\frac{\log\lambda_{j}}{\log r}\in\mathbb{Q}$ for $j\in J$. Let $j\in J$ and pick for $n\geq 1$ a positive integer $k_{n}$ so that
@@
For any $1\leq t\leq p$, let $q_{t,n}$ be the largest positive integer such that $r^{q_{t,n}}\leq\delta^{n}/a_{t}$. Since $r_{i}=r$ for all $i\in I$, $I^{\prime}_{t,n}=I^{q_{t,n}}$ (recall (4.3)). Then for $t\neq t^{\prime}$, we always have $a_{t}r^{q_{t,n}}\leq\delta^{n}<a_{t^{\prime}}r^{q_{t^{\prime},n}-1}$ and hence $r^{q_{t^{\prime},n}-q_{t,n}}>a_{t}r/a_{t^{\prime}}$. So there is a constant $C\geq 1$ such that
(4.6) $|q_{t,n}-q_{t^{\prime},n}|\leq C,\quad\forall 1\leq t,t^{\prime}\leq p,\forall n\in\mathbb{Z}^{+}.$
By Lemma 4.3 and Proposition 4.4,
(4.7) $\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\Big{\{}\mathcal{H}^{s}(K)\sum_{t=1}^{p}c_{n,t}a_{t}^{s}r^{q_{t,n}s}:c_{n,t}\in\mathbb{Z}^{+}\cup\{0\}\text{ with }\sum_{t=1}^{p}c_{n,t}\leq P\Big{\}}.$
Writing
$\mathcal{B}:=\Big{\{}\sum_{t=1}^{p}c_{t}a_{t}^{s}r^{\beta_{t}s}:c_{t}\in\mathbb{Z}^{+}\cup\{0\},\beta_{t}\in\mathbb{Z}\text{ are such that}\sum_{t=1}^{p}c_{t}\leq P\text{ and}-C\leq\beta_{t}\leq C\Big{\}},$
we see from (4.6) and (4.7) that
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{j^{k_{n}}}))\in\{\mathcal{H}^{s}(K)r^{q_{1,n}s}b:b\in\mathcal{B}\},\quad\forall n\geq 1.$
Since $\mathcal{B}$ is a finite collection, we can find (by the pigeonhole principle) a pair of distinct integers $n,\tilde{n}$ such that
@@
Unfortunately, the above argument fails without the homogeneity of $\Phi$, and even granting the homogeneity of $\Psi$ seems to offer little help. This reveals a subtle asymmetry: to upgrade a Lipschitz embedding to a Lipschitz isomorphism, assuming the homogeneity on $K$ is markedly easier than assuming that on $F$. It is noteworthy that Proposition 4.4 may relate to a characterization given by Xi [20] via finding appropriate graph-directed subsystems. However, we do not know whether the measure-preserving property in Proposition 4.4 would directly ensure the condition in Xi’s criterion (and thereby conclude the equivalence).
@@
    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))\approx\mathcal{H}^{s}(F_{\bm{j}})$ for all $1\leq t\leq p$ and $\bm{i}\in\mathcal{I}_{t}$;
@@
    $\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0$ unless $(t,\bm{i})=(t^{\prime},\bm{i}^{\prime})$.
@@
The first statement is simple: since $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\subset F_{\bm{j}}$, the $\lesssim$ direction holds trivially; for the inverse direction, one can apply Remark 4.5 to $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ and get
$\geq\widetilde{c}^{-1}\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}}+b_{t})\approx\mathcal{H}^{s}(F_{\bm{j}}),$
where the last estimate holds because $\lambda_{\bm{j}}\approx a_{t}r_{\bm{i}}$ when $\bm{i}\in I^{\prime}_{t,n}$.
It remains to verify the second statement. Since $\widetilde{f}:\widetilde{K}\to F$ is a surjection,
$F_{\bm{j}}=\widetilde{f}\big{(}\widetilde{f}^{-1}(F_{\bm{j}})\big{)}=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}).$
Let $(t,\bm{i})\neq(t^{\prime},\bm{i}^{\prime})$. Write $E:=\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}})$ and suppose on the contrary that $\mathcal{H}^{s}(E)>0$. Then $E$ is a Borel subset of the dust-like self-similar set $F_{\bm{j}}$ with positive $\mathcal{H}^{s}$-measure. By [6, Lemma 2.12], we can find positive integers $m_{1}<m_{2}<\cdots$, and $\eta_{k}\in J_{m_{k}}$ (with $F_{\eta_{k}}\subset F_{\bm{j}}$) such that
(4.8) $\mathcal{H}^{s}(E\cap F_{\eta_{k}})>\big{(}1-k^{-1}\big{)}\mathcal{H}^{s}(F_{\eta_{k}}),\quad\forall k\geq 1.$
For each $k\geq 1$, consider $A_{k}:=\widetilde{f}^{-1}(F_{\eta_{k}})$ and let
@@
Clearly, $A_{k,1}\cap A_{k,2}=\varnothing$ (since $(t,\bm{i})\neq(t^{\prime},\bm{i}^{\prime})$) and $A_{k,1}\cup A_{k,2}\subset A_{k}$. Note that
(4.9) $\widetilde{f}(A_{k,1})=F_{\eta_{k}}\cap\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$
(because $\widetilde{f}$ is onto and thus $\widetilde{f}(\widetilde{f}^{-1}(A)\cap B)=A\cap\widetilde{f}(B)$ for all $A\subset F$ and $B\subset K$). Combining this with $\mathcal{H}^{s}(E\cap F_{\eta_{k}})>0$, we see that $A_{k,1}\neq\varnothing$. Similarly, $A_{k,2}\neq\varnothing$. Moreover, it follows from (4.8) and (4.9) that
@@
By Lemma 4.3, we can write $\widetilde{f}^{-1}(F_{\eta_{k}})=\bigcup_{\ell=1}^{p}\bigcup_{\omega\in\mathcal{I}_{\ell,m_{k}}}(a_{\ell}O_{\ell}K_{\omega}+b_{\ell})$ for some $\mathcal{I}_{\ell,m_{k}}\subset I^{\prime}_{\ell,m_{k}}$. Since $F_{\eta_{k}}\subset F_{\bm{j}}$,
$A_{k,1}=\bigcup_{\omega\in\mathcal{I}_{t,m_{k}}:K_{\omega}\subset K_{\bm{i}}}(a_{t}O_{t}K_{\omega}+b_{t})\quad\text{and}\quad A_{k,2}=\bigcup_{\omega\in\mathcal{I}_{t^{\prime},m_{k}}:K_{\omega}\subset K_{\bm{i}^{\prime}}}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}}).$
By Proposition 4.4, $\psi_{\eta_{k}}^{-1}\widetilde{f}^{-1}(F_{\eta_{k}})\in\mathcal{K}$. So $\psi_{\eta_{k}}^{-1}(A_{k,1}),\psi_{\eta_{k}}^{-1}(A_{k,2})\in\mathcal{K}$ (because they are non-empty sub-unions). Also
$\mathcal{H}^{s}(A_{k,2})$ $\geq\min_{\omega\in I^{\prime}_{t^{\prime},m_{k}}}\mathcal{H}^{s}(a_{t^{\prime}}O_{t^{\prime}}K_{\omega}+b_{t^{\prime}})$
$\geq(\underline{r}\delta^{m_{k}})^{s}\mathcal{H}^{s}(K)$ (recall (4.3))
$\geq\frac{\underline{r}^{s}\Delta_{F}^{s}}{L^{s}}\cdot\lambda_{\eta_{k}}^{s}\mathcal{H}^{s}(K)$ (since $\eta_{k}\in J_{m_{k}}$)
@@
$\psi_{\eta_{k}}^{-1}(x)$ $\mapsto\psi_{\eta_{k}}^{-1}\widetilde{f}(x)$
with $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$. Since $\psi_{\eta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$, we can find by Lemma 2.2 a set $B\in\mathcal{K}$ and a Lipschitz map $g:B\to F$ with $\operatorname{Lip}(g)\leq L$ such that
$\mathcal{H}^{s}(B)\leq\sup_{k}\mathcal{H}^{s}(\psi_{\eta_{k}}^{-1}(A_{k,1}))=\mathcal{H}^{s}(F)\cdot\sup_{k}\frac{\mathcal{H}^{s}(A_{k,1})}{\mathcal{H}^{s}(F_{\eta_{k}})}<\widetilde{c}\mathcal{H}^{s}(F)$
@@
$\mathcal{H}^{s}(g(B))$ $\geq\limsup_{k\to\infty}\mathcal{H}^{s}(g_{k}(\psi_{\eta_{k}}^{-1}(A_{k,1})))$
@@
The latter implies that $g(B)=F$, i.e., $g:B\to F$ is a Lipschitz surjection with $\operatorname{Lip}(g)\leq L$. However, it follows from $B\in\mathcal{K}$ and our choice of $\widetilde{K}$ (recall Corollary 4.2) that
$\widetilde{c}\mathcal{H}^{s}(F)>\mathcal{H}^{s}(B)\geq\mathcal{H}^{s}(\widetilde{K})=\widetilde{c}\mathcal{H}^{s}(F),$
which is a contradiction. This proves (1). ∎